# FRAP analysis Measuring biophysical kinetic parameters using image analysis

### v1

@@ III Results and discussion
Given the application to FRAP experiments, one challenge facing our nuclear segmentation and tracking algorithms is generating masks for the bleached nucleus, which often vanishes from both channels post-bleaching. To overcome this challenge, we used the relative positions of neighboring nuclei and assumed the bleached nucleus moved in a coordinated fashion with the surrounding nuclei and that it maintained its shape (Fig. 10a). As such, nucleus tracking also affects the segmentation of the bleached nucleus in the frames during which it has vanished. To test these assumptions, we conducted a comparative analysis between masks generated under normal conditions and those derived using the neighboring location-based method, employing Intersection over Union (IoU) as our evaluation metric. The comparative results indicate that our approach yields an average IoU of .85 across more than 100 nuclei and 18 embryos, suggesting a high degree of accuracy (Fig. 10b-f). This performance metric further improves when the size of the nucleus remains consistent over the observed period.
@@
Our work contributes an effective pipeline for segmenting and tracking nuclei, tailored specifically for biological data where such processes are inherently challenging yet fundamentally important. This pipeline proves especially valuable in contexts requiring accurate identification and continuous monitoring of nuclei to decipher cellular behaviors and interactions. The enhanced precision and efficiency in tracking provided by our pipeline offer substantial support to researchers engaged in examining complex biological environments. It facilitates a deeper understanding of nuclear dynamics, which is pivotal for advancing scientific knowledge in fields reliant on detailed cellular analysis.
The implications of our study extend beyond the immediate results, highlighting the transformative potential of our pipeline in changing the way researchers conduct segmentation and tracking of nuclei across diverse biological settings to extract quantitative measurements of biophysical parameters.

### v3

@@ III-B Tracking
Given the application to FRAP experiments, one challenge facing our nuclear segmentation and tracking algorithms is generating masks for the bleached nucleus, which sometimes vanishes from both channels post-bleaching. We used the approach described in the Materials and Methods section to overcome this challenge. As such, nucleus tracking also affects the segmentation of the bleached nucleus in the frames during which it has vanished (Fig. 11a). To test these assumptions, we conducted a comparative analysis between masks generated under normal conditions and those derived using the neighboring location-based method, employing Intersection over Union (IoU) as our evaluation metric. The comparative results indicate that our approach yields an average IoU of .85 across more than 100 nuclei and 18 embryos, suggesting a high degree of accuracy (Fig. 11b-f). This performance metric further improves when the size of the nucleus remains consistent over the observed period.
III-C Extracting intensities of nuclei and cytoplasm
The nucleus tracking algorithm described in the Materials and Methods section allowed us to accurately and consistently identify the fluorescence recovery of the bleached nucleus (Fig. 9a). Furthermore, our approach also extracts the intensities of the surrounding nuclei and cytoplasmic regions (Fig. 9a). These intensities of the surrounding nuclei and cytoplasm (Fig. 9a) show a slight dip which would indicate the occurrence of unintentional bleaching. We normalized the intensity of the bleached nucleus and its associated cytoplasm (Fig 9b) using 10 of its closest neighbors. The resulting normalized intensities (Fig. 9b) ensure that the measurements reflect true biological variations rather than artifacts of the imaging process.
III-D Estimating import and export rates
The final analytical phase involved using the normalized fluorescence intensity data (Fig. 9b) to estimate protein import and export rates. By fitting these intensities using the procedure described in the Methods section, we determined the rates at which the protein entered and exited the nucleus. Given the biological variation among the embryos, the boxplot of the import and export rates for Dl at two locations (dorsal and ventral) are presented (Fig. 10). The results show the variation in the rates based on the location which indicates that the spatial asymmetry of the Dl concentration.
@@ III-D Estimating import and export rates
Our work contributes an effective pipeline for segmenting and tracking nuclei, tailored specifically for biological data where the measurement of biophysical parameters from experimental image data is inherently challenging yet fundamentally important. This pipeline proves valuable in contexts requiring accurate identification and continuous monitoring of nuclei to decipher cellular behaviors and interactions.